The ball dropped at 7m building has a velocity of 2.5m/s, it reaches the ground at




10.2s? What is the distance travelled by the ball?




What are the given:





What is asked in the problem:




Formula to use:




Solution:

1
Expert's answer
2022-06-14T13:39:12-0400

Given:

"h=7\:\rm m"

"v=2.5\:\rm m/s"

"t=10.2\:\rm s"

Unknown:

"d"



"d=gt^2/2\\
=9.8*10.2^2/2=510\:\rm m"